OpenStudy (anonymous):

Can someone find the derivative of y=ln (cos(x^2+4)/cos(x^2+9))^5

OpenStudy (raden):

u need the division rule, also the chain rule

OpenStudy (anonymous):

I know, I just made it up to practice and I need someone else to work on it so I know I get the right answer.

OpenStudy (raden):

have u tried that ?

OpenStudy (anonymous):

I'm working on it right now. I just need someone else to work on it too so I have something to check my answer with

OpenStudy (raden):

Oh, great... :) i afraid my answer will wrong :P

OpenStudy (anonymous):

i'm still working on my answer

OpenStudy (anonymous):

I'm having a hard time putting my answer into the equation option on here, post your answers and I will work on posting mine

OpenStudy (anonymous):

\[5\ln \frac{(\cos(x^2+4)) }{(\cos(x^2+9)) }\frac{ (-2xsin(x^2+4)\cos(x^2+9))-(-2xsin(x^2+9)\cos(x^2+4)) }{(\cos(x^2+4)(\cos(x^2+9)) ? }\]

OpenStudy (anonymous):

@RadEn and @Hoa I posted my answer

OpenStudy (anonymous):

I guess mine is wrong? @jim_thompson5910 @Directrix can you come and give the answer to this question?

OpenStudy (raden):

wait. is that like :